OpenStudy (anonymous):

is this right? y+5=-12 y=7

OpenStudy (ranga):

No. To get y subtract 5 from both sides.

OpenStudy (anonymous):

7 + 5 = 12 What would make that Y + 5 = -12? The answer is -17 + 5 = 12 :)

OpenStudy (anonymous):

urg..i knew i got it wrong! It was too easy lol.

OpenStudy (anonymous):

Haha It's ok I get my Negative and Positive's mixed up all the time just think about money and the number line :) -10 -9 -8 -7 -6 -5 -4 -3 -2 -1 0 1 2 3 4 5 6 7 8 9 10 If you have -7 and Add 2 you go Forward not backwards :) Because you are adding to a NEGATIVE number :) If you have 1 and Subtract 5 just go backwards and im sure You can get it now? :D

OpenStudy (anonymous):

well if you just have one and you go backwards wouldn't that just be zero? So if you had -3 + 2 you would go backwards?

OpenStudy (anonymous):

oh you said subtract 5. sorry it cut off

OpenStudy (anonymous):

No it would go forward :) =-1

OpenStudy (anonymous):

If a negative and a negative are close like this 1 - -1 = 2 You add :)

OpenStudy (anonymous):

Oh okay, you might have said this already but when you add to a negative number you go forward and when you subtract from negative it goes back and if you add to a positive you go back? and when you add you go forward? Or is it the other way around for positive?

OpenStudy (anonymous):

+ and + go Forward :) - and - go forward :) - and + go forward + and - go back :) I think that's about it :)

OpenStudy (anonymous):

Oh my haha

OpenStudy (anonymous):

I think I understand though! Thank you haha!

OpenStudy (anonymous):

It's easy once you get the hang of it I promise!! :)

OpenStudy (anonymous):

Hope I helped! :D Thankyou!

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

Why are you thanking me' ? haha